If a current of 0.5 ampere flows througha mettalic wire for  2 hours , then how many electrons would flow through the wire ?

Grade:12

1 Answers

Robin Chourasia
10 years ago
USING Q=It

=> Q= 3600 coulombs

No. of e= Q/charge of e

N= 3600/(1.6 * 10^-19) = 2.2 *10^22